FEL Upgrade Project Monthly Report April 2002

This month we were pleased to welcome a visit by Virginia's new Secretary of Technology, George Newstrom and the Deputy Secretary, Eugene Huang on April 4th.  Our guests were given a tour of CEBAF and the FEL Facility and participated in a ceremony honoring Jefferson Lab's recent patent award winners.

We are pleased to acknowledge receipt of a valuable piece of equipment that will benefit the UV FEL Upgrade program.  Brookhaven National Lab has shipped to us a prototype version of the Undulator that was built by a Cornell University-Argonne National Lab collaboration.  The undulator was most recently carefully characterized during use in the Brookhaven’s " High Gain High Harmonic Experiment".  This equipment is being loaned to the JLab FEL Program on a long term loan arrangement.  We thank our colleagues at Cornell and Brookhaven for the valuable loan and useful characterization data.

The first girder assemblies for the e-beam transport system were assembled during April.  Preliminary testing started this month on the 300 kW High Voltage Power Supplies (HVPS) for the Injector klystrons.

Considerable attention was devoted this month by both FEL project management and the project team for preparations for the project semiannual review which was held on May 1-3.  A component of the preparations for the review included a careful analysis of the remaining budget for completion of the 10 kW upgrade project and the commissioning activities.

We were pleased to receive the funding from ONR for the first segment ($1.3M) of recently approved FY02 FEL project activities in the amended ONR/DOE Memorandum of Agreement, including support of commissioning , the fabrication of the 3rd cryomodule and studies on short wavelength wigglers.

A revised "Work for Others" contract for the FY02 increment in funding for the UV FEL project was forwarded to our AFRL contract monitor from the DOE JLab Site Office.  This contract revision would add an additional $1.5M to the work scope based on FY02 DOD Appropriations for the JLab/Aerospace collaboration on the UV FEL project.

An article was published this week in the May/June issue of the "Industrial Physicist" that summarizes materials science applications of the FEL and the upcoming "Industrial Physics Forum" which will be hosted by JLab in October 2002.  The article can be found on the Web at:

Gwyn Williams attended a DOE sponsored workshop this month in Charlotte on the "Genomes to Life" program which included discussions of potential applications of the planned synchrotron addition to the FEL Facility.

On a less positive note, we found out this month that the anticipated funding for the $3M addition to FEL facility to house the Helios synchrotron did not make the FY03/04 Virginia state budget approved on April 17th by the General Assembly.  We have been told by the Governor's Office that the administration intends to restore the budget item during the next General Assembly session in January 2003.
Project Cost Performance:
The Phase 1 FEL Upgrade project budget for the period June 1, 2000 to September 30, 2001 is $9,029k.  The first phase of the project was fully obligated as of the end of September.  The total Phase 1 project cost to date-actual costs plus commitments is $9,029k as planned for Phase 1 completion.

The Phase 2 FEL Upgrade project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002.  Actual funds transferred from ONR to DOE for this effort in FY02 were $4,442k.  Work for Phase 2 proceeded to plan during this reporting period.  Actual charges of $317k were accrued for April for total accrued costs of $3018k to date for Phase 2.  Work scheduled for April was $436k for a total of $2,109k for Phase 2 to date; work performed for April was $317k for a total of $2,438k for Phase 2 to date, which results in 55% scheduled vs. 47% performed for the Phase 2 project.  At present the cost variance for Phase 2 is -$580k and the schedule variance is $329k.  We expect to meet the planned cost at completion of Phase 2 effort.  The program is presently on track to begin commissioning efforts in September 2002.

The UV FEL project was approved for $2,836k for a performance period from September 2001 to September 30, 2002.  Actual charges of $183k were accrued through April for a total accrued costs of $831k since project start.  Work scheduled for April was $95k for a total of $955k for the UV project to date, work performed for April was $170k for a total of $1,048k for the UV project to date, which results in 37% scheduled vs. 34% performed for the UV project.  At present the cost variance for the UV project is $217k and the schedule variance is $93k.  The program is on track to meet the statement of work requirements on schedule and within the available budget.

WBS 3 (Beam Physics):
We are still slogging through field quality related issues for dipoles and quads.  Documentation of dipoles is in progress and discussions of quad powering are underway (EPICS asynchronicity makes this entertaining).  Thanks to Jay Benesch for reopening this can of worms and Mike Tiefenback for sharing his considerable experience with CEBAF (sort of a "by his stripes we are healed" kind of thing).

Dipole field characterization review suggests dipoles (with shaving) can provide operationally acceptable core field and  central field integrals at energies we'll be able to reach for the near future.  AES is characterizing behavior with shims - we'll see what 210 MeV/c looks like soon.

Moving on to field homogeniety, we've projected the beam stay-clear onto AES results - figures modified from Tom Schultheiss's results follow - these show the field variation across the beam is pretty much at the 10^-4 level we need.

Cogitation on commissioning continues - perusal of IR Demo log entries from commissioning suggests a six month effort is reasonable provided an adequate number of victimizable participants are available to run 2 shifts/day, 5 days/week (order 10 experienced accelerator physicist/engineering staff).

Discussion of space (or lack thereof) for the multi-slit diagnostic immediately upstream of the linac led Benson to make the very amusing observation that the device, as presently installed, is not on a bellows and thus may not be very well aligned.  This could go a long way towards explaining why it’s hard to fit beam through it.  Were Jock still here, he'd certainly award Steve this week's Blue Star award for that call!

WBS 4 (Injector):
Continued working to support the W&M implantation system.  A new support/mount for the support tube was made along with 2 tubes for test runs.  Polishing on the support tube and ball cathode continued.  Detail part drawings for the new resistor are being worked.

The gun stand and the extractor cart have been delivered this month along with the 5" diameter ceramic cylinders for the Conditioning Resistor.

Minor modifications to the BPM/ bellows assembly after the lightbox was resolved.  The BPM will be manufactured and tested.

Gun HVPS Tank - The last of the large transmission line piping  and the old HVPS tank was removed this month in preparation for the new tank and transmission line slated for delivery by early June.

PARMELA simulations are underway regarding fine-tuning of the telescope quads to optimize e-beam parameters at the entrance of the first LINAC.

WBS 6 (RF):
Quarter HVPS - HV wiring in HVPS #3 is nearly complete, except for wiring to the klystron cart.  Changed out all meters for HVPS #3, except for 2 that have not yet been received.  The control wiring for HVPS #3 is about 80% complete.  The large capacitors were secured to the chassis of HVPS #3 and the missing MOV's were located. Work continued on HVPS # 4.  The control wiring is about 60% complete.  More meters were received and installed this week.  A different PLC is being ordered to simplify the communication task with EPICS.  A major portion of this communication software is already written.

Quarter Klystrons - Plumbing work on the final 2 klystron carts is expected to start soon by the Machine Shop.

WBS 8 (Instrumentation):
The final cleanup off the cable trays took place this month.  Many thanks to the I&C students and Jim!  For the upcoming new cable installation, the beam viewers will be followed by the trim magnets.  Both of these systems have the new documentation complete.  The software screens are nearly complete for the viewers, this was due to assistance from the CEBAF operations crews; thanks.

The fast raster power supplies upgrade has begun.  There are minor modifications to the drive and control printed circuit board that need to be done.  The markups are nearly complete, they will then enter the EECAD que.  The design for the upgraded high voltage power supply control is also proceeding.  This task will combine three chassis into one and present a much "cleaner" interface to the Machine Protection System (MPS).

The design and documentation for the Beam Loss Monitor VME conversion is nearly complete, and a prototype PCB is being fabricated.  Front and rear panels are already in fabrication.

As a part of the cable tray cleanup the damaged ion pump HV cables were removed, we are awaiting the response from the cable manufacturer on resolving the cable failures.  To prevent any electrical hazards the remaining high voltage power supplies were relocated to the tunnel so that NO active high voltage cables remain in the tray.

The fabrication of the individual beam viewer and camera interface boxes is nearly complete.  The interface boxes provide the interface to the control system, the Machine Protection System (MPS) and power for the cameras and lamp rings.  Seven (7) out of 36 new 3" shielded viewers (for both IR & UV machines) remain to be wired.  During the construction of the last FEL this work was done post installation in the accelerator enclosure, by doing it in the lab there is ample time for the students to do the wiring, and same time during commissioning.  Each of these are fully tested and signed off then sent on to the vacuum group for girder installation.  The few viewers that remained installed will be refitted with the new interface as time allows.  The new software screens for viewer and video control are nearly complete.

The design is complete for the compact 1.75" injector BPM.  The new design has a slot length of ~3" which should allow easy installation with bolts (not studs) and enable the use of a standard bellows upstream.  The channel assignments for Switched Electrode Electronics (SEE) and 4 channel electronics have also been made.  The software changes will be done in a few weeks.

The beam viewer control chassis were installed into the racks at FL03B09, the debugging of the software screens and field termination will begin next week.  The system print for the beam viewer system has also been started.  This is one of the more strait forward tasks to be done so we're knocking it off first.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
At Magnet Enterprises International (MEI) in Oakland CA.  The third coil looks very good.  The fourth is being prepared for potting , the fifth coil looks very good.  The 6th is in potting.
Return legs remain in final machining.  Top and bottom slabs are in grinding and 8 (for 4 magnets) are promised as a batch for this month.  Then the machining subcontractor will do final machining on the batch to create four magnet cores.
Injector Dipoles (DU/DV)
• WANG NMR wound all the GU and GV Coils.  They are being ground wrapped in fiberglass tape in prep for final potting.
• Master Machine is nearly finished rough machining the slabs for the yokes.  Heat treating will be performed soon. Covers are complete. Field clamps are finished.  Mu metal sheets for the GU did not come in to the uniform thickness specified and we are working with the supplier to correct that.
Arc Dipoles (GY, GX, GQ)
• The contract for GQ-GX magnet cores was established with Process Equipment Co. of the Dayton Ohio area.  They did all the yokes for the original CEBAF accelerator. The steel for GQ-GX magnet cores was shipped to Process Equipment Co.  AES is magnetically modeling the GQ-GX dipoles at lower energy with shorter pole tips to match David Douglas’ use matrix.  They will then design a shim system as part of Revision A for the pole tips to reconfigure the magnets for high energy use.
• The contract for the coils and assembly of the GQ-GX was established with Wang NMR.  They are doing a very competent job at winding GY, GG, GU and GV coils.  The copper for the GQ-GX coils was received by Wang NMR.  Wang NMR is ready to wind the GY coil as soon as they wrap the insulation around the conductor. They are making the potting mold for the GY coils.  They were delayed in starting impregnation of the GG coils by late delivery of epoxy.
• Bosma Machine has the GY core pieces in final planing of the pole surface and the return legs in finish machining before grinding.
3 inch quad (QX)
• Some questions arose about the production measurements and hysteresis cycles matching the cycles in actual use and about the stability of the power supply being used.  At month’s end, these questions are near resolution.  In any case, there are no quadrupoles that are outside the specification for multipole content.  Their slight differences in strength will be taken care of by individualized curves in the control system.
• We decided to use the full bipolar voltage (“Bang-bang”) method of running through hysteresis loops advocated by Mike Tiefenback, magnet integrator for CEBAF.  Our production measurements have not used this hysteresis cycle and we are testing several magnets to determine if there is a constant offset to the data that can be applied to our existing readings.
• The power supply being used to test the quads was observed to have an unacceptable ripple.  At month’s end, due to these questions, we have developed a circuit to smooth the ripple.  We are also making sure the power supply has no regulation problems.  We then plan to continue the measurements.
• Eight more of the UV quads were received.
Trim Quad (QT)
• Milhous Control of Virginia reports that potting a first article was delayed.  Potting didn't take place because a sample potting attempt caught a bad batch of epoxy that would not polymerize.  A replacement batch is in transit.
Sextupole (SF)
• DULY Research performed additional optimization aimed at simplifying construction.  At the same time, a construction method concept using quadrupole core machining technology is being incorporated in the mechanical design.  Our thanks to New England Technicoil for discussions that led to this concept.

Octupole (OT)
• Work on this magnet is on furlough until the sextupole is designed.
Beam Line and Vacuum
• Design of the ARC Chambers at AES continues, incorporating the new 100 kW aperture in the second arc GQ chambers.  Overall design is very mature, with some chamber detail drawings at their second stage of review.
• The assembly and detail drawings for the girder for the region at the end of the cryomodules was signed.
• The assembly and detail drawings of the girders for the six quadrupole telescope in front of the first arc was completed, checked and is in final review at last.
• The Region in front of the first cryomodules including the injection line is laid out and detailed, ready for first review.  The region includes stands and the X Chamber for the Injection dipoles.

WBS 10 (Wiggler):
We took final measurements on the wiggler this month.  The first measurements showed that the wigglers had the wrong polarity for X-ray work.  The leads were reversed and the field checked.  The corrector strengths were the same as before to within measurement precision so changing the polarity has no effect on the wiggler calibration.  The interaction of the wiggler and dispersion section lead to a reduction in the strength of the last half pole in the wiggler as expected and a slightly reduction in the outer pole of the dispersion section (not expected but a small effect).  The net result of this is that the wiggler corrector coils will have to run about 1.5% stronger than previous calibrations indicated.  Otherwise no problems seen.  Took some measurements of the noise offsets and started staging for vacuum chamber reinstallation.  We moved the wiggler to the vault on April 24.

WBS 11 (Optics):
The leadscrew and nut hardware are being procured.  Design details continue on the vacuum vessels.  Modeling of the 50kW laser beam dump is essentially complete for one concept (the flow channel design).  Some results, such as the likelihood and impact of a decentered beam, need to be discussed and resolved.  Design of the chamber and optics for the THz light pickoff continue.

It will be installed in the OCR, and tested once the LSOP is modified and approved.  The drive laser compressor hardware is installed in the drive laser enclosure and is being aligned before testing begins.
Bids are being solicited for the new vacuum valves for the optical cavity chambers and optical transport system.  We received the bids on the optical cavity internals and made the award.  Detailing of the lead screw holding assembly is nearly complete and should be completed next week.  As the designer for the vacuum vessels is unavailable, that effort was reassigned.  We held a meeting to bring the new designer up to speed and are making good progress. Design of the chamber and optics for the THz light pickoff continue.  We are evaluating using an ultraviewer form the IR Demo to actuate the pickoff mirror.

The pump laser for the ultrafast laser system arrived, and we moved the laser system into UL 1, where it will be initially checked out, once the LSOP is modified and approved.  We have discovered that the lens that couples the drive laser output into the single-mode fiber is creating diffraction rings because of the diamond tooling process.  This is unacceptable for our pulse compressor, so we are looking at other vendors.  Work continues on the OCMMS piezo transducer controls.  The mechanical mount has resonance at ~ 50, 110, and 150 Hz, and so the feedback circuit frequency response has to be lowered to avoid them.  We are adding a transducer to the mirror mount to mimic low frequency building vibrations, and then watching how well the circuit counteracts them.  The OCMMS mirrors are being procured.  We are soliciting a new round of estimates for the other cavity optics and the transport optics.

We held a kick-off meeting on the DIRUP-funded (through the College of William and Mary) work we will do to support these valuable additions to the facility.